OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride? 27:8 9:6 9:4 3:2

OpenStudy (anonymous):

@jim_thompson5910 can you please help me?

jimthompson5910 (jim_thompson5910):

The two cones are similar. So we know that r1/r2 = k for some constant k In this case, r1 = 33 and r2 = 22, so r1/r2 = k 33/22 = k 3/2 = k k = 3/2 This means that we're multiplying the smaller radius (22 ft) by 3/2 or 1.5 to get the larger radius (33 ft) Since the two cones are similar, the larger height (h1) is found by multiplying the smaller height (h2) by 1.5 So h2 = 1.5*h1 Now turn to the formula of a cone V = (pi*r^2*h)/3 and then plug in r = 33 and h = x to get V = (pi*33^2*x)/3 V = (1089pi*x)/3 V = 363pi*x So the volume of the larger cone is exactly 363pi*x cubic feet. Now let's plug in r = 22 and h = y to find the volume of the smaller cone V = (pi*r^2*h)/3 V = (pi*22^2*y)/3 V = (484pi*y)/3 So the volume of the smaller cone is exactly (484pi*y)/3 cubic feet. Since the larger height is 1.5 times the smaller height, we can say x = 1.5y which means 363pi*x turns into 363pi*1.5y 544.5pi*y So the volume of the larger cone is also exactly 544.5pi*y cubic feet Now divide the two volumes (larger into smaller) (Larger Volume)/(Smaller Volume) (544.5pi*y)/((484pi*y)/3) 3.375 Since 3.375 converts to the improper fraction 27/8, this means that the ratio is 27:8

jimthompson5910 (jim_thompson5910):

That seems like a lot of work...but...it leads you to the very useful shortcut if the ratio of two corresponding lengths is a:b, then the ratio of the volumes is a^3:b^3

OpenStudy (anonymous):

wow thanks and you actually explained it to where i understand. thanks @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

That's great. I'm glad I could help.